sonu's grandmother is 80 years old and sonu is 20 year old . how many years ago was his grandmother 7 times as old as sonu ?
Answer by addingup(3677)

Today grandma is 80 and Sonu is 20. This is smoke screen information, you don't need it.
:
Here is what we have to find:
find when
80-x = 7(20-x)
80-x = 140-7x add 7x and subtract 80 on both sides
6x = 60
x = 10 10 years ago is our answer
-------------------------
Check:
10 years ago, grandma was:
80-10 = 70
and Sonu was:
20-10 = 10
Grandma was 7 times older than Sonu, so Sonu times 7 = grandma:
10*7 = 70 Correct